OpenStudy (anonymous):

Please help algebra 2

OpenStudy (anonymous):

Tony is trying to find the equation of a quadratic that has a focus of (−1, 4) and a directrix of y = 8. Describe to Tony your preferred method for deriving the equation. Make sure you use Tony's situation as a model to help him understand.

OpenStudy (anonymous):

first start with the vertex-form of a parabola, its: y = a(x-h)^2 + k next, solve for the vertex of the quadratic by using the value of the focus and the directrix. since, the focus is: (-1, 4) and the direxrix is: y=8 use the mid point formula from the focus to the directrix to find the y-value of the vertex. the x-value of the vertex will be the same as the x-value in the focus, -1 so at this point, we will have the values of the vertex, (h, k). plug those into the equation. all that is left to find is the value for 'a'

OpenStudy (anonymous):

Would the midpoint be -0.5/ 2?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

hmm.... maybe it would be easier if I showed it this way:  where is the midpoint for this number line, given the two points, i circle?

OpenStudy (anonymous):

the midpoint would be 6

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

so we know that the vertex of the parabola is (h, k) or (-1, 6) plug those into the equation: y = a(x-h)^2 + k

OpenStudy (anonymous):

y = a(-1 - 6)^2 + k?

OpenStudy (anonymous):

close!! (h, k) is (-1, 6) so, h = -1 k = 6 try again ^_^ once more

OpenStudy (anonymous):

y = a(x- (-1))^2 + 6

OpenStudy (anonymous):

yep, thats right

OpenStudy (anonymous):

so, two more steps, the next step is to figure out what 'a' is. then once we have that, we expand the equation we have been 'making'

OpenStudy (anonymous):

Could you keep explaining how to solve? I know I may be frustrating you but I really do want a good grade in my class.

OpenStudy (anonymous):

yea, i can do that

OpenStudy (anonymous):

do you have any question about why i've done the steps I have so far?

OpenStudy (anonymous):

Nope, I think I've gotten it. I'm writing it down on a sheet of paper and following along :)

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

so the next step was to find 'a' 'a' is the distance from the vertex (-1, 6) to the focus (-1, 4) can you tell me what that distance is?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

 try again ^_^

OpenStudy (anonymous):

Would it be -1, 5? or just 5? I used the distance formual to get 12 lol

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

did u get my message i sent? a=-2